# On complexity of restricted fragments of Decision DNNF

### v1

@@ 4 The effect of structuredness on Decision dnnf
An important property of obdd is efficiency of conjunction of two obdds obeying the same order. In particular, if $B_{1}$ and $B_{2}$ are two obdds over the same set of variables and obeying the same order $\pi$ then the function $f(B_{1})\wedge f(B_{2})$ can be represented by an obdd obeying $\pi$ of size $|B_{1}|\cdot|B_{2}|$ and can be computed at time $O(|B_{1}|\cdot|B_{2}|)$. This nice property is lost by fbdd. In fact, it is not hard to demonstrate that if two obdd $B_{1}$ and $B_{2}$ obey different orders then the size an fbdd representing $f(B_{1})\wedge f(B_{2})$ can be exponentially lower-bounded by $|B_{1}|+|B_{2}|$.
dnnf, being a generalization of fbdd, does not enable efficient conjunction. A restriction on dnnf that enables efficient conjunction is structuredness. The purpose of this section is to define two different forms of structuredness for Decision dnnfs and discuss the power of these models.
@@
Let $D$ be a dnnf A $\wedge_{d}$ node $u$ of $D$ with inputs $u_{1}$ and $u_{2}$ respects a vtree $VT$ if there is a node $x$ of $VT$ with children $x_{1}$ and $x_{2}$ such that $\mathsf{var}(D_{u_{1}})\subseteq\mathsf{var}(VT_{x_{1}})$ and $\mathsf{var}(D_{u_{2}})\subseteq\mathsf{var}(VT_{x_{2}})$. A dnnf $D$ over $V$ respects $VT$ if each node of $D$ respects $VT$ A dnnf $D$ is structured if it respects a vtree.
To apply the structuredness to Decision dnnf, we need to be aware of two equivalent definitions of the model: as a special case of dnnf and as a generalization of fbdd. So far, we have used the latter definition which we referred to as $\wedge_{d}$-fbdd. The former definition states that a Decision dnnf is a dnnf where each $\vee$ node $u$ obeys the following restriction. Let $u_{0}$ and $u_{1}$ be the inputs of $u$. Then they both are $\wedge_{d}$ nodes and there is a variable $x\in\mathsf{var}(D)$ such that one of inputs of $u_{0}$ is an input gate labelled with $\neg x$ and one of the inputs of $u_{1}$ is an input gate labelled with $x$. Then a structured decision dnnf is just a $\wedge_{d}$-fbdd whose representation as a Decision dnnf respects some vtree. As the fpt upper bound of [11], in fact, applies to structured Decision dnnfs, the following holds.
@@
Yet the model is quite restrictive and can be easily ’fooled’ into being xp-sized even with one long clause. In particular, the following has been established in [3].
@@
Let $G$ be a graph without isolated vertices. Let $\varphi^{*}(G)$ be a cnf obtained from $\varphi(G)$ by adding a single clause $\neg V$. Then the size of structured Decision dnnf representing $\varphi^{*}(G)$ is at least $\Omega(\textsc{fbdd}(G))$.
In particular, there is an infinite set of positive integers $k$ for each of them there is a class ${\bf G}_{k}$ of graphs of treewidth at most $k$ (and hence the $itw(\varphi^{*}(G))\leq k+1$) and the structured Decision dnnf representation being of size at least. $n^{\Omega(k)}$ where $n=|V(G)|$.
In order to get an additional insight into the ability of $\wedge_{d}$-dnnf to efficiently represent cnfs of bounded incidence treewidth, we introduce an alternative definition of structuredness, this time based on $\wedge_{d}$-fbdd.
@@
A $\wedge_{d}$-fbdd $B$ respects a vtree $VT$ if each $\wedge_{d}$ node of $B$ respects $VT$. That is, unlike in the case of the ’proper’ structuredness, we do not impose any restrictions on the decision nodes of $B$. We say that $B$ is structured if $B$ respects a vtree $VT$.
A structured $\wedge_{d}$ fbdd is a generalization of fbdd as it does not impose any restriction on a $\wedge_{d}$-fbdd in the absence of $\wedge_{d}$ nodes. As a result the property of efficient conjunction held for structured dnnf does not hold structured $\wedge_{d}$-fbdds.
@@
Note that, even though Decision dnnf and $\wedge_{d}$-fbdd, their structured versions are not.
@@
We note that $\psi[{\bf g}]$ is not necessarily the same function as $\psi|_{\bf g}$ as the set set of the variables of the former may be a strict subset of the set of variables of the latter. Yet, it is not hard to demonstrate the following.
@@
In order to prove Theorem 18, we identify a set ${\bf C}$ of at most $p$ ’long’ clauses whose removal from $\varphi$ results in a cnf of primal treewidth at most $k$. An important part of the resulting structured $\wedge_{d}$-fbdd is representation of ${\bf C}$ by an $O(n^{p})$-sized decision tree. We are now going to define a decision tree and to demonstrate the possibility of such a representation.
@@
A decision tree is a rooted binary tree where each leaf node is labeled by either $true$ or $false$ and each non-leaf node is labelled with a variable. Moreover, for each non-leaf node one outgoing edge is labelled with $1$, the other with 00. A decision tree must be read-once, that is the same variable does not label two nodes along a directed path. Then each directed path $P$ corresponds to the assignment ${\bf a}(P)$ in the same way as we defined it for $\wedge_{d}$-fbdd.
For a decision tree $DT$, $\mathsf{var}(DT)$ is the set of variables labelinbg the non-leaf nodes of $DT$. We say that $DT$ represents a function $f$ is $\mathsf{var}(DT)\subseteq\mathsf{var}(f)$ and the following two statements hold.
@@
If $\varphi$ has no clauses then the resulting decision tree consists of a single $true$ node. If $\varphi$ has an empty clause then the decision tree consists of a single $false$ node.
Otherwise, pick a clause $C$ with variables $x_{1},\dots,x_{r}$. We define $DT(\varphi)$ as having a root-leaf path $P$ consisting of vertices $u_{1},\dots,u_{r+1}$ where $u_{1},\dots,u_{r}$ are respectively labelled with $x_{1},\dots,x_{r}$. Furthermore for each $1\leq i\leq r$, the outgoing edge of $u_{i}$ is labelled with 00 if $x_{i}\in C$ and with $1$ if $\neg x_{i}\in C$. Clearly, ${\bf a}(P)$ falsifies $C$, hence we label $u_{r+1}$ with $false$.
Next, for each $1\leq i\leq r$, we introduce a new neighbour $w_{i}$ of $v_{i}$ and, of course, label the edge $(u_{i},w_{i})$ with the assignment of $x_{i}$ as in $C$. Let $P_{i}$ be the path from the root to $w_{i}$ and let ${\bf g}_{i}={\bf g}(P_{i})$ Clearly, ${\bf g}_{i}$ satisfies $C$. Therefore $\varphi[{{\bf g}_{i}}]$ can be expressed as a cnf of at most $p-1$ clauses. Make each $w_{i}$ the root of $DT(\varphi[{{\bf g}_{i}}])$. Clearly, the resulting tree represents $\varphi$. By the induction assumption, the trees $DT(\varphi[{{\bf g}_{i}}])$ are of size at most $n^{p-1}$ and there are at most $|C|\leq n$ such trees. Hence the lemma follows. ∎
@@
    Let $B^{\prime}$ be the dag obtained as a result of the above two stages. Let $X=\mathsf{var}(\varphi)\setminus\mathsf{var}(B^{\prime})$. If $X=\emptyset$ we set $B=B^{\prime}$. Otherwise, let $x_{1},\dots,x_{a}$ be the variables of $X$. Then $B$ is obtained from $B^{\prime}$ by introducing new nodes $u_{1},\dots,u_{a}$ and setting $u_{a+1}$ to be the source of $B^{\prime}$. For each $1\leq i\leq a$, label $u_{i}$ with $x_{i}$ and introduce two edges from $u_{i}$ to $u_{i+1}$, one labelled with $1$, the oth er labelled with 00.
@@ 5 Conclusion
The main purpose of this paper is obtaining better understanding of the power of $\wedge_{d}$-fbdd in terms of the representation of cnfs of bounded incidence treewidth.
For that purpose, we considered several restrictions on $\wedge_{d}$. The first considered model is $\wedge_{d}$-obdd. The xp lower bound for $\wedge_{d}$-obdd representing cnfs of bounded incidence treewidth leads us to the first open question. One motivation for considering $\wedge_{d}$-fbdd for cnf of bounded incidence treewidth is that this would provide an insight into the understanding of complexity of regular resolution for such cnfs. The xp lower bound for $\wedge_{d}$-obbd thus naturally raises a question whether such a lower bound would hold for the corresponding restriction over the regular resolution.
It is well known that a regular resolution refutation of a cnf $\varphi$ can be represented as a read-once branching program whose sinks are labelled with clauses of $\varphi$. The constraint is that the assignment carried by each source-sink path must falsify the clause associated with the sink. Let us say that the a regular resolution refutation is oblivious if the corresponding read-once branching program obeys a linear order of its variables.
@@
What is the complexity of oblivious regular resolution parameterized by the incidence treewidth?
One important property of obdds is a possibility of efficient conjunction of two obdds obeying the same order: such obdds can be seen as structured dnnfs respecting the same vtree. One important consequence is a polynomial time algorithm for satisfiability testing of conjunction of two obdds obeying the same order. To put it differently, the conjunction of two obdds obeying the same order meets the minimal requirement for being considered a knowledge compilation model. It is thus interesting whether the same holds for conjunction of two $\wedge_{d}$-obdds obeying the same order.
@@
Let $B_{1}$ and $B_{2}$ be two $\wedge_{d}$ obdds obeying the same order $\pi$. Is there a polynomial time algorithm for testing whether $f(B_{1})\wedge f(B_{2})$ has at least one satisfying assignment?
Finally, we believe that understanding the complexity of structured $\wedge_{d}$-fbdd representing cnfs of bounded incidence treewidth is an important step towards understanding such complexity for $\wedge_{d}$-fbdds. In this context, we will pose a more specific open question related to Theorem 18.
@@
The upper bound presented in Theorem 18 is fpt in $k$ but xp in $p$. Does an fpt upper bound parameterized by $k+p$ hold for structured $\wedge_{d}$-fbdd representing the class of cnfs as specified in Theorem 18?

### v2

@@ 5 The effect of structuredness on Decision dnnf
dnnf, being a generalization of fbdd, does not enable efficient $Apply$ (conjunction) operation. A restriction on dnnf that enables carrying $Apply$ efficiently is structuredness [16]. The purpose of this section is to define two different forms of structuredness for Decision dnnfs and discuss the power of these models.
@@
Let $D$ be a dnnf A $\wedge_{d}$-node $u$ of $D$ with inputs $u_{1}$ and $u_{2}$ respects a vtree $VT$ if there is a node $x$ of $VT$ with children $x_{1}$ and $x_{2}$ such that $\mathsf{var}(D_{u_{1}})\subseteq\mathsf{var}(VT_{x_{1}})$ and $\mathsf{var}(D_{u_{2}})\subseteq\mathsf{var}(VT_{x_{2}})$. A dnnf $D$ over $V$ respects $VT$ if each $\wedge_{d}$-node of $D$ respects $VT$. A dnnf $D$ is structured if it respects a vtree.
To apply the structuredness to Decision dnnf, we need to be aware of two equivalent definitions of the model: as a special case of dnnf and as a generalization of fbdd. So far, we have used the latter definition which we referred to as $\wedge_{d}$-fbdd. The former definition states that a Decision dnnf is a dnnf where each $\vee$ node $u$ obeys the following restriction. Let $u_{0}$ and $u_{1}$ be the inputs of $u$. Then they both are $\wedge_{d}$-nodes and there is a variable $x\in\mathsf{var}(D)$ such that one of inputs of $u_{0}$ is an input gate labelled with $\neg x$ and one of the inputs of $u_{1}$ is an input gate labelled with $x$. Then a structured decision dnnf is just a $\wedge_{d}$-fbdd whose representation as a Decision dnnf respects some vtree. As the fpt upper bound of [14], in fact, applies to structured Decision dnnfs, the following holds.
@@
Yet the model is quite restrictive and can be easily ‘fooled’ into being xp-sized even with one long clause. In particular, the following has been established in [5].
@@
Let $G$ be a graph without isolated vertices. Let $\varphi^{*}(G)$ be the cnf obtained from $\varphi(G)$ by adding a single clause $\neg V(G)=\{\neg v|v\in V(G)\}$. Then the size of structured Decision dnnf representing $\varphi^{*}(G)$ is at least $\Omega(\textsc{fbdd}(\varphi(G)))$.
In particular, there is an infinite set of positive integers $k$ for each of them there is a class ${\bf G}_{k}$ of graphs of treewidth at most $k$ (and hence the $itw(\varphi^{*}(G))\leq k+1$) and the structured Decision dnnf representation of $\varphi^{*}(G)$ being of size at least. $n^{\Omega(k)}$ where $n=|V(G)|$.
In order to get an additional insight into the ability of $\wedge_{d}$-fbdd to efficiently represent cnfs of bounded incidence treewidth, we introduce an alternative definition of structuredness, this time based on $\wedge_{d}$-fbdd.
@@
A $\wedge_{d}$-fbdd $B$ respects a vtree $VT$ if each $\wedge_{d}$-node of $B$ respects $VT$. That is, unlike in the case of the ‘proper’ structuredness, we do not impose any restrictions on the decision nodes of $B$. We say that $B$ is structured if $B$ respects a vtree $VT$.
@@
Note that, even though Decision dnnf and $\wedge_{d}$-fbdd are the same model, their structured versions are not. In particular, a structured $\wedge_{d}$-fbdd is a generalization of fbdd as it does not impose any restriction on a $\wedge_{d}$-fbdd in the absence of $\wedge_{d}$ nodes. As a result the property of efficient conjunction held for structured dnnf does not hold for structured $\wedge_{d}$-fbdds.
@@
We note that $\psi[{\bf g}]$ is not necessarily the same function as $\psi|_{\bf g}$ as the set of the variables of the former may be a strict subset of the set of variables of the latter. Yet, it is not hard to demonstrate the following.
@@
In order to prove Theorem 34, we identify a set $\bf C$} of at most $p$ ’long’ clauses whose removal from $\varphi$ results in a cnf of primal treewidth at most $k$. An important part of the resulting structured $\wedge_{d}$-fbdd is an $O(n^{p})$-sized decision tree representation of $\bf C$}. We are now going to define a decision tree and to demonstrate the possibility of such a representation.
@@
A decision tree is a rooted binary tree where each leaf node is labeled by either $True$ or $False$ and each non-leaf node is labelled with a variable. Moreover, for each non-leaf node one outgoing edge is labelled with $1$, the other with 0. A decision tree must be read-once, that is the same variable does not label two nodes along a directed path. Then each directed path $P$ corresponds to the assignment ${\bf a}(P)$ in the same way as we defined it for $\wedge_{d}$-fbdd.
For a decision tree $DT$, $\mathsf{var}(DT)$ is the set of variables labeling the non-leaf nodes of $DT$. We say that $DT$ represents a function $f$ is $\mathsf{var}(DT)\subseteq\mathsf{var}(f)$ and the following two statements hold.
@@
If $\varphi$ has no clauses then the resulting decision tree consists of a single truetrue node. If $\varphi$ has an empty clause then the decision tree consists of a single falsefalse node.
Otherwise, pick a clause $C$ with variables $x_{1},\dots,x_{r}$. We define $DT(\varphi)$ as having a root-leaf path $P$ consisting of vertices $u_{1},\dots,u_{r+1}$ where $u_{1},\dots,u_{r}$ are respectively labelled with $x_{1},\dots,x_{r}$. Furthermore for each $1\leq i\leq r$, the outgoing edge of $u_{i}$ is labelled with 0 if $x_{i}\in C$ and with $1$ if $\neg x_{i}\in C$. Clearly, ${\bf a}(P)$ falsifies $C$, hence we label $u_{r+1}$ with $False$.
Next, for each $1\leq i\leq r$, we introduce a new neighbour $w_{i}$ of $u_{i}$ and, of course, label the edge $(u_{i},w_{i})$ with the assignment of $x_{i}$ as in $C$. Let $P_{i}$ be the path from the root to $w_{i}$ and let ${\bf g}_{i}={\bf g}(P_{i})$ Clearly, ${\bf g}_{i}$ satisfies $C$. Therefore $\varphi[{{\bf g}_{i}}]$ can be expressed as a cnf of at most p1p-1 clauses. Make each $w_{i}$ the root of $DT(\varphi[{{\bf g}_{i}}])$. Clearly, the resulting tree represents $\varphi$. By the induction assumption, the trees $DT(\varphi[{{\bf g}_{i}}])$ are of size at most $n^{p-1}$ and there are at most $|C|\leq n$ such trees. Hence the lemma follows. ∎
@@
    Let $B^{\prime}$ be the dag obtained as a result of the above two stages. Let $X=\mathsf{var}(\varphi)\setminus\mathsf{var}(B^{\prime})$. If $X=\emptyset$ we set $B=B^{\prime}$. Otherwise, let $x_{1},\dots,x_{a}$ be the variables of $X$. Then $B$ is obtained from $B^{\prime}$ by introducing new nodes $u_{1},\dots,u_{a}$ and setting $u_{a+1}$ to be the source of $B^{\prime}$. For each $1\leq i\leq a$, label $u_{i}$ with $x_{i}$ and introduce two edges from $u_{i}$ to $u_{i+1}$, one labelled with $1$, the oth er labelled with 0.
@@ 6 Conclusion
The main purpose of this paper is obtaining better understanding of the power of $\wedge_{d}$-fbdd in terms of the representation of cnfs of bounded incidence treewidth. We believe that the next important step forward is understanding the complexity of structured $\wedge_{d}$-fbdd representing cnfs of bounded incidence treewidth. Recall that Theorem 34 present an upper bound that is fpt in $k$ but xp in $p$. A natural question for further consideration is whether we can improve the upper bound so that it is fpt. We conjecture that this is not the case, a formal statement of the conjecture is provided below.
Conjecture 1.
There are constants $c$ and $d$ such that the following holds. For each sufficiently large $k$, there is an infinite class ${\bf\Phi}_{k}$ of cnfs $\varphi$ so that $\varphi$ can be turned into a cnf of primal treewidth $ck$ by removal of $k$ clauses and such that a representation of $\varphi$ as a structured $\wedge_{d}$-fbdd is of size at least $n^{d\cdot k}$ where $n=|Var(\varphi)|$.
We believe a recent result [21] will be useful for resolving Conjecture 1. In [21], we consider $\wedge_{d}$-fbdd with imbalanced gates. Roughly speaking, each $\wedge_{d}$-gate has only one input depending on more than $n^{\alpha}$ inputs for some constant $0<\alpha<1$ fixed for a specific $\wedge_{d}$-fbdd. We establish a lower bound $n^{\Omega(1-\alpha)\cdot k}$ for cnf of primal treewidth. Intuitively, this means that Conjecture 1 can be considered only for balanced vtrees. In particular, we believe that if an xp lower bound is obtained for balanced vtrees the lower bound can be combined with the lower bound in [21] to form a ‘win-win’ argument.
We believe that the xp lower bound for $\wedge_{d}$-obdd has potential to be ‘transferred to the area of proof complexity. Indeed, a similar ‘primal vs incidence treewidth’ dilemma is known for the resolution proof system: it is easy to demonstrate existence of fpt sized resolution refutation parameterized by the primal treewidth of the input cnf. However, the existence of fpt sized resolution refutation parameterized by the incidence treewidth is a well known open question. Importantly, the same dilemma persists for the regular resolution.
It is well known that a regular resolution refutation of a cnf $\varphi$ can be represented as a read-once branching program whose sinks are labelled with clauses of $\varphi$. The constraint is that the assignment carried by each source-sink path must falsify the clause associated with the sink. Let us say that the a regular resolution refutation is oblivious if the corresponding read-once branching program respects a linear order of its variables. It is known that although being similar to fbdd, regular resolution has a greater intuitive similarity to $\wedge_{d}$-fbdd: in order to refute a conjunction of two variable-disjoint cnfs, it is enough to refute just one of them. Continuing this similarity to the oblivious case, we observe an intuitive similarity between $\wedge_{d}$-obdd and oblivious regular resolution. The xp lower bound for the former naturally leads to an open question whether a similar lower bound exists for the latter.
@@
What is the complexity of oblivious resolution refutation parmeterized by the incidence treewidth of the input cnf? In particular, can an xp lower bound be established?
In the rest of the section, we will consider possible directions of further research stemming from our results related to the Apply operation. Recall, that we demonstrated a possibility of efficient Apply operation for two $\wedge_{d}$-obdd embeddable into the same order if they are both ‘close’ to an obdd respecting the same order. The obdd can be seen as a ‘common denominator’ of two $\wedge_{d}$-obdds. A natural direction is to consider closeness of two models to each other rather than to some third model. This leads us to stating the following ‘open-ended’ open question.
@@
Introduce a parameter measuring closeness of two $\wedge_{d}$-obdds $B_{1}$ and $B_{2}$ respecting the same order but not necessarily embeddable into it so that $f(B_{1})\wedge f(B_{2})$ can be represented as a $\wedge_{d}$-obdd of size $|B_{1}|\times|B_{2}|$ to the power proportional to the value of the parameter.
Question 3 naturally leads to a similar question being asked about structured dnnf.
@@
Introduce a parameter $k=k(T_{1},T_{2})$ measuring closeness of two vtrees $T_{1}$ and $T_{2}$ so that if $B_{1}$ and $B_{2}$ are two structured dnnfs respecting $T_{1}$ and $T_{2}$ respectively then $f(B_{1})\wedge f(B_{2})$ can be represented as a structured dnnf of size $(|B_{1}|\cdot|B_{2}|)^{\Theta(k)}$.
It is known that the notion of embeddability, introduced in this paper in the context of $\wedge_{d}$-obdds is applicable to dnnfs and even to Boolean circuits in general [1]. This leads us to the following open question consisting of two parts.
Open question 5.
  1. 1.
    Introduce the notion of the irregularity index of a dnnf embeddable into a specific linear order.
  2. 2.
    A nondeterministic read-once branching program is oblivious if there is a linear order of variables, so that along each source-sink path, the variables are queried according to this order. We refer to such a branching program as a Nondeterministic obdd (nobdd) (the model is discussed in Section 10.1 of [25]). Let $B$ be a dnnf embeddable into an order $\pi$ with irregularity index $k$. Similarly to Theorem 19, can $B$ be simulated by a nobdd of size $|B|^{O(k)}$?
It is known that the embeddability can be seen as a weaker form of structuredness [1]. The goal of the next conjecture is to make this connection ‘quantified’.
Conjecture 2.
A structured dnnf is also a dnnf embeddable into a specific order with the irregularity index at most $\log n$.
Specifically, we believe that, for a vtree $T$ a linear order $\pi$ of variables and the irregularity index $k$ of $T$ w.r.t. this order can be naturally defined so that $k\leq\log n$ where $n$ is the number of variables of $T$. Furthermore, we believe that if $B$ is a structured dnnf respecting $T$ then $B$ is embeddable into $\pi$ with irregularity index at most $k$.
Conjecture 2 needs to be accompanied with demonstration of succinctness of embeddable dnnfs as compared to structured ones. This connection is formally specified in the next two open questions. The first question seeks to establish succinctness of embeddable dnnfs as compared to structured ones. The second question is a refinement of the first one subject to a specific irregularity index.
Open question 6.
Is there an infinite set $\bf F$} of Boolean functions so that each $f\in{\bf F}$ can be represented by an embeddable dnnf of size polynomial in $|\mathsf{var}(f)|$ but requires a structured dnnf of size exponential in $|\mathsf{var}(f)|$ for its representation?
Open question 7.
Does the following hold for each sufficiently large $k$. There is an infinite class ${\bf F}_{k}$ of Boolean functions so that each $f\in{\bf F}_{k}$ is representable as an embeddable dnnf with irregularity index $k$ and of size polynimial in $|\mathsf{var}(f)|$ but requires for its representation an exponential sized structured dnnf respecting a vtree with irregularity index at most $k$.
The final open question that we propose in this paper facilitates the primary purpose of the embeddability notion, namely efficient implementation of Apply. The open question below can be seen as an upgrade of Open Question 4.
Open question 8.
Introduce a parameter $k=k(B_{1},B_{2})$ measuring closeness of two embeddable dnnfs $B_{1}$ and $B_{2}$ so that $f(B_{1})\wedge f(B_{2})$ can be represented as an embeddable dnnf of size at most $(|B_{1}|\cdot|B_{2}|)^{\Theta(k)}$. In case $B_{1}$ and $B_{2}$ are structured respecting trees $T_{1}$ and $T_{2}$, respectively, the parameter should be equal $k(T_{1},T_{2})$ as in Open Question 4.